# SPAR: Self-supervised Placement-Aware Representation Learning for Distributed Sensing

### v3

@@ 4.3 Evaluation on RealWorld-HAR Dataset
Method RealWorld-HAR Recognition Accuracy (%) () F1 (%) () SPAR 89.63 0.57 90.45 0.63 SPAR+LLM 90.40 0.68 91.13 0.59
@@
We also evaluate SPAR+LLM, elaborated in Section 3.1, which replaces learnable structural positions with LLM-derived embeddings from textual sensor metadata. On this dataset, where we can create rich natural language descriptions of sensor placement, SPAR+LLM yields additional gains over SPAR (Table 4). This highlights the benefit of combining spatial priors with semantic context for placement-aware learning.
@@ 5 Conclusion
This paper presents SPAR, a general self-supervised pretraining framework designed for the whole spectrum of multi-modal, multi-node distributed sensing. By introducing spatial and structural positional embeddings alongside dual reconstruction objectives, SPAR leverages the inherent duality between observer positions and observed signals to enable placement-aware representation learning. Theoretical analyses grounded in information theory and occlusion-invariant learning offer principled support for the framework. Extensive experiments across three real-world datasets, spanning diverse sensing modalities, placement configurations, and task types, demonstrate the superior generalizability and robustness of SPAR. We hope SPAR inspires broader efforts toward integrating spatial and structural context into foundational representation learning paradigms for the wide range of distributed sensing applications.
Ethics Statement
This work develops a general pretraining framework for distributed sensing using only publicly available datasets that contain no personally identifiable or sensitive information. Our study does not involve human subjects or private data collection, and we believe it poses no direct ethical concerns.
Reproducibility Statement
We have taken concrete steps to ensure the reproducibility of our results. Detailed descriptions of datasets, preprocessing procedures, and training protocols are provided in Appendix E. Formal proofs of our theoretical results are included in Appendix D. In addition, we release our implementation and scripts at https://anonymous.4open.science/r/SPAR-4C74/.
@@ Appendix A Additional Experiments Results
Method M3N-VC Single-vehicle Localization Message Drop Rate 0.05 Message Drop Rate 0.1 Message Drop Rate 0.2 MSE () () Dist. Err. () () MSE () () Dist. Err. () () MSE () () Dist. Err. () () CMC 50.13 17.69 6.48 0.91 50.27 18.03 6.60 1.17 55.38 22.01 6.91 1.29 Cosmo 29.97 2.30 5.44 0.12 29.37 2.43 5.44 0.10 33.19 2.67 5.61 0.12 SimCLR 26.77 2.55 5.16 0.07 25.65 1.28 5.15 0.07 28.31 3.94 5.4 0.18 AudioMAE 19.29 1.42 4.91 0.17 18.25 1.23 4.77 0.10 19.03 1.14 4.77 0.05 CAV-MAE 16.28 0.17 4.68 0.03 15.85 0.81 4.57 0.10 15.98 1.67 4.44 0.12 FOCAL 26.62 1.02 5.21 0.17 27.42 2.33 5.32 0.20 33.03 2.28 5.70 0.15 FreqMAE 27.48 1.43 5.14 0.11 28.32 1.22 5.17 0.14 28.68 6.96 5.32 0.24 PhyMask 23.18 4.96 4.98 0.32 24.09 3.31 5.03 0.27 27.37 2.04 5.31 0.21 SPAR 12.65 0.61 4.09 0.11 12.48 0.68 4.07 0.02 14.56 2.47 4.27 0.07
Unseen Sensor Placements. Next, we assess generalization to unseen sensor placements. All models are pretrained on the full dataset excluding scenes H08 and H24 (which share similar configurations), then finetuned and evaluated on H24. To simulate transfer, nodes in H24 are assigned structural position vectors randomly drawn from those learned during pretraining. As reported in Table 7, SPAR continues to outperform baselines, underscoring its placement-aware generalization ability.
Method M3N-VC Single-vehicle Localization (Finetuned and Evaluated on Unseen Placement) MSE () () Dist. Err. () () CMC 61.78 17.68 7.23 0.78 Cosmo 63.43 12.85 7.22 0.63 SimCLR 35.82 7.57 5.92 0.39 AudioMAE 41.25 4.60 6.64 0.31 CAV-MAE 37.01 0.68 6.27 0.04 FOCAL 41.79 11.04 5.91 0.41 FreqMAE 30.65 1.14 5.51 0.19 PhyMask 34.83 8.81 5.60 0.38 SPAR 21.76 1.00 5.09 0.10
Ablations. Finally, we conduct ablations to quantify the contribution of each design choice (Table 8). Removing geometric augmentation, the spatial reconstruction objective, or spatial embeddings all leads to comparable performance drops, highlighting their complementary roles. Excluding structural embeddings results in the most severe degradation, underscoring their critical role in modeling node-specific characteristics. We also test alternative masking strategies: Node-Drop Masking (masking entire nodes) reduces performance, while Node-Balanced Masking (ensuring a minimum number of unmasked tokens per node) offers slight gains over random masking. We further vary the mask ratio (0.85 and 0.5) and observe only minor performance changes relative to the default 0.75, indicating the robustness of the framework.
Method M3N-VC Single-vehicle Localization MSE () () Dist. Err. () () SPAR 12.98 0.11 4.20 0.07 w/o Geometric Augmentation in Pretrain 15.59 0.56 4.67 0.04 w/o Reconstructing Spatial Positions 14.73 0.35 4.62 0.03 w/o Spatial Positional Embedding 15.12 0.58 4.67 0.07 w/o Structural Positional Embedding 22.55 2.98 5.08 0.13 + Node-Drop Masking 17.71 3.17 4.82 0.33 + Node-Balanced Masking 12.54 1.63 4.14 0.21 + Mask Ratio 0.85 13.89 2.17 4.35 0.24 + Mask Ratio 0.5 14.81 2.57 4.45 0.31
@@ Appendix D Proofs
In this proof we use and to denote generic constants independent of model parameters, whose specific values may change from Equation to Equation.
Classical MAE. We begin with the case of classical MAE. We assume, following prior works (Li et al., 2022b; Kong & Zhang, 2023), that due to the high dimension of the latent embeddings relative to the original signal, the latent embeddings may contain the full information about the unmasked part of the signals , which can be reconstructed by the decoder from the latent embeddings with negligible loss. As a result, we consider the reconstruction loss calculated on the masked signals to be equivalent to the reconstruction loss calculated on the full signals:
@@
SPAR. For SPAR, the signal decoder takes additional inputs: masked spatial and structural positional embeddings (Equation 3). Let and denote the masked spatial and structural positions. Let denote the signal reconstruction loss for modality . Then, adjusting our reasoning above, we can modify Equation 13 to:
@@
Classical MAE. Kong et al. (Kong & Zhang, 2023) provided a rigorous interpretation of classical MAE as a special case of contrastive learning, where the positive pair consists of two complementary masked views of the same input signals. For completeness and clarity, we briefly restate their reasoning here using our notation. For clarity, we focus on a single modality by omitting the superscript and the joint encoder ; the extension to multiple modalities is straightforward.
@@
Kong et al. (Kong & Zhang, 2023) assumes that due to the high dimension of the latent embeddings relative to the original signals, the latent embeddings produced by the may approximately preserve all the information of the input. This implies the existence of an alternative decoder that can satisfy:
@@
Following the same argument of Kong et al. (Kong & Zhang, 2023), this shows that in SPAR can be viewed as a contrastive loss between two masked views of the signal, enriched with shared spatial and structural context:
@@ E.1 Baseline Methods Descriptions
CMC (Tian et al., 2020) Learns shared representations by maximizing mutual information between views, enabling view-agnostic and scalable contrastive learning across multiple modalities.
Cosmo (Ouyang et al., 2022) Integrates contrastive feature alignment with attention-based selective fusion to effectively capture shared and distinctive patterns from multimodal data under scarce labeling.
SimCLR (Chen et al., 2020) Forms discriminative visual embeddings by aligning augmented image pairs through a nonlinear projection and optimizing the NT-Xent contrastive loss.
AudioMAE (Huang et al., 2022) Applies masked autoencoding to audio by operating on spectrogram patches, using a Transformer to reconstruct masked regions and capture time-frequency patterns without relying on external modalities.
CAV-MAE (Gong et al., 2022) Combines masked autoencoding and contrastive learning in a unified audio-visual framework, using modality-specific encoders and a joint decoder to learn both fused and aligned representations from spectrogram and image patches.
FOCAL (Liu et al., 2023) Separates multimodal time-series signals into shared and private latent spaces, enforcing orthogonality and applying contrastive and temporal constraints to capture both modality-consistent and modality-exclusive features.
FreqMAE (Kara et al., 2024b) Enhances masked autoencoding for multimodal sensing by incorporating frequency-aware transformers, factorized fusion of shared and private features, and a weighted loss that prioritizes informative frequency bands and high-SNR samples.
PhyMask (Kara et al., 2024a) Improves masked autoencoding by adaptively selecting time-frequency patches based on energy and coherence metrics, enabling efficient and informative masking tailored to physical sensing signals.
@@ E.2 Settings for Multi-modal Multi-node Vehicle Classification Dataset
The Multi-Modality Multi-Node Vehicle Classification Dataset (M3N-VC) (Li et al., 2025) (CC BY 4.0) comprises synchronized audio and vibration recordings of four vehicle types, along with background noise, collected from March 2023 to October 2024. Data were gathered across six distinct real-world scenes, each featuring diverse terrain types (asphalt, dirt, gravel, and concrete) and varying weather conditions (sunny, rainy, and windy).
Each scene is instrumented with a spatially distributed sensor network composed of 6 to 8 nodes. Every node includes a co-located microphone (sampled at 16 kHz) and a geophone (sampled at 200 Hz). Vehicle GPS trajectories were recorded at a rate of 1 Hz. All recordings are segmented into non-overlapping 2-second clips, resulting in a total of 21,694 samples. These clips are transformed into mel-scale spectrograms for model input. GPS coordinates are converted into meter-level spatial positions using the Local Tangent Plane approximation (Agency, 1987).
@@
Multi-Vehicle Joint Classification and Localization. For multi-vehicle settings, we use scene I22, which contains multiple moving vehicles. A two-layer transformer serves as the task head, trained with a DETR-style loss function (Carion et al., 2020) to handle set-based predictions.
@@ E.3 Settings for Ridgecrest Seismicity Dataset
This dataset contains seismic waveform recordings from 31,452 earthquake events (M > -0.5) occurring between January 1, 2020, and December 31, 2024, within an 80 km radius of (35.9°, -117.6°) in the Ridgecrest region of California. The data collection and processing procedures largely follow the methodology outlined by Si et al. (Si et al., 2024).
We obtained the earthquake event catalog by querying the Southern California Seismic Network (SCSN) (California Institute of Technology (1926), Caltech) via the Southern California Earthquake Data Center (SCEDC) (Center, 2013) online catalog. The selected events include magnitudes higher than -0.5 and depths larger than than -5 km. For each event, we collected three-component (East, North, Vertical) waveform data from 16 stations in the California Integrated Seismic Network, using two modalities: high-gain broadband seismometers and high-gain accelerometers. All data are sampled at 100 Hz and retrieved in miniSEED format from the SCEDC Open Data repository.
For each event, we extract a 30.72-second window from all channels as model input. During preprocessing, we detrend the waveforms and apply the Short-Time Fourier Transform (STFT) to generate spectrograms. A 2-35-Hz band-pass filter is applied to remove low-frequency noise (e.g., oceanic and atmospheric microseisms) and high-frequency instrumental or environmental noise. We convert spatial positions of each station from GPS signals to kilo-meter-level positions using Local Tangent Plane projection (Agency, 1987).
@@ E.4 Settings for RealWorld-HAR Dataset
The RealWorld Human Activity Recognition (HAR) dataset (Sztyler & Stuckenschmidt, 2016) comprises multi-modal activity signals collected from 15 participants. The dataset captures eight common activity types: walking, sitting, lying, climbing down, running, standing, climbing up, and jumping.
@@
Appendix F LLM Usage Statement
LLMs were used in a supportive role for polishing the writing and providing occasional coding assistance, with all outputs carefully verified by the authors. Technical ideas, experimental designs, and theoretical analyses were developed by the authors. The authors take full responsibility for the final content of this paper.

### v4

@@ 5 Conclusion
This paper presents SPAR, a self-supervised pretraining framework designed for the full spectrum of multi-modal, multi-node distributed sensing. By introducing spatial and structural positional embeddings alongside dual reconstruction objectives, SPAR leverages the inherent duality between observer positions and observed signals to enable placement-aware representation learning. Theoretical analyses grounded in information theory and occlusion-invariant learning offer principled support for the framework. Extensive experiments across three real-world datasets, spanning diverse sensing modalities, placement configurations, and task types, demonstrate the superior generalizability and robustness of SPAR. We hope SPAR inspires broader efforts toward integrating spatial and structural context into foundational representation learning paradigms for the wide range of distributed sensing applications.
Impact Statement
This work advances self-supervised representation learning for distributed sensing systems by explicitly incorporating sensor placement into pretraining. By enabling more robust, generalizable, and data-efficient learning across diverse sensing modalities and layouts, the proposed framework has the potential to improve real-world applications such as environmental monitoring, infrastructure safety, transportation systems, and human activity analysis. More accurate and resilient sensing models can support better decision-making, resource allocation, and situational awareness in both civilian and scientific contexts. This research focuses on methodological and foundational advances in representation learning, and we do not anticipate direct negative societal impacts arising from the proposed approach.
@@ Appendix A Additional Experiments Results
Preservation of Spatial Information in Learned Representations. We conduct an additional ablation study to directly examine whether SPAR preserves spatial information in its learned representations. Using the M3N-VC dataset under the single-vehicle localization setting, we randomly perturb each node’s spatial position by adding i.i.d. uniformly distributed noise with varying magnitudes. The pretrained encoder is frozen, and a lightweight spatial decoder is trained to reconstruct the perturbed spatial positions from the learned representations. We compare SPAR with an ablated variant that does not include the spatial reconstruction objective during pretraining. As shown in Figure 3(a), the reconstruction error for the ablated model increases approximately linearly with the magnitude of the injected noise, while SPAR maintains consistently low reconstruction error across all noise levels, providing direct evidence that SPAR’s representations robustly encode spatial position information.
Robustness to Noise in Spatial Positions. We further investigate the robustness of SPAR to noise in sensor spatial positions by perturbing the position of a single node with uniformly distributed random noise in the M3N-VC single-vehicle localization task. We consider two settings: injecting noise only at test time, and injecting noise consistently during both fine-tuning and testing. Figure 3(b) shows that localization error increases as the magnitude of spatial perturbation grows, indicating that the downstream localization head actively relies on spatial information encoded in the representations. At the same time, performance remains stable under small perturbations, deteriorates smoothly when the noise increases, and stays reasonable even when the maximal perturbation magnitude exceeds the average distance to the nearest neighboring node. These observations suggest that SPAR’s representations capture spatial context without depending on precise coordinates of individual nodes, demonstrating the robustness to realistic spatial noises.
@@ Appendix D Proofs
In this proof, we use and to denote generic constants independent of model parameters, whose specific values may change from Equation to Equation.
Classical MAE. We begin with the case of classical MAE. We assume, following prior works (li2022exploring; kong2023understanding), that due to the high dimension of the latent embeddings relative to the original signal, the latent embeddings may contain the full information about the unmasked part of the signals , which can be reconstructed by the decoder from the latent embeddings with negligible loss. As a result, we consider the reconstruction loss calculated on the masked signals to be equivalent to the reconstruction loss calculated on the full signals:
@@
SPAR. For SPAR, the signal decoder takes additional inputs: masked spatial and structural positional embeddings. Let and denote the masked spatial and structural positions. Let denote the signal reconstruction loss for modality . Then, adjusting our reasoning above, we can modify Equation 3 to:
@@
Classical MAE. Kong et al. (kong2023understanding) provided a rigorous interpretation of classical MAE as a special case of contrastive learning, where the positive pair consists of two complementary masked views of the same input signals. For completeness and clarity, we briefly restate their reasoning here using our notation. For clarity, we focus on a single modality by omitting the superscript and the joint encoder ; the extension to multiple modalities is straightforward.
@@
Kong et al. (kong2023understanding) assumes that due to the high dimension of the latent embeddings relative to the original signals, the latent embeddings produced by the may approximately preserve all the information of the input. This implies the existence of an alternative decoder that can satisfy:
@@
Following the same argument of Kong et al. (kong2023understanding), this shows that in SPAR can be viewed as a contrastive loss between two masked views of the signal, enriched with shared spatial and structural context:
@@ E.1 Baseline Methods Descriptions
CMC (tian2020contrastive) Learns shared representations by maximizing mutual information between views, enabling view-agnostic and scalable contrastive learning across multiple modalities.
Cosmo (ouyang2022cosmo) Integrates contrastive feature alignment with attention-based selective fusion to effectively capture shared and distinctive patterns from multimodal data under scarce labeling.
SimCLR (chen2020simple) Forms discriminative visual embeddings by aligning augmented image pairs through a nonlinear projection and optimizing the NT-Xent contrastive loss.
AudioMAE (huang2022masked) Applies masked autoencoding to audio by operating on spectrogram patches, using a Transformer to reconstruct masked regions and capture time-frequency patterns without relying on external modalities.
CAV-MAE (gong2022contrastive) Combines masked autoencoding and contrastive learning in a unified audio-visual framework, using modality-specific encoders and a joint decoder to learn both fused and aligned representations from spectrogram and image patches.
FOCAL (liu2023focal) Separates multimodal time-series signals into shared and private latent spaces, enforcing orthogonality and applying contrastive and temporal constraints to capture both modality-consistent and modality-exclusive features.
FreqMAE (kara2024freqmae) Enhances masked autoencoding for multimodal sensing by incorporating frequency-aware transformers, factorized fusion of shared and private features, and a weighted loss that prioritizes informative frequency bands and high-SNR samples.
PhyMask (kara2024phymask) Improves masked autoencoding by adaptively selecting time-frequency patches based on energy and coherence metrics, enabling efficient and informative masking tailored to physical sensing signals.
@@ E.2 Settings for Multi-modal Multi-node Vehicle Classification Dataset
The Multi-Modality Multi-Node Vehicle Classification Dataset (M3N-VC) (li2025restoreml) (CC BY 4.0) comprises synchronized audio and vibration recordings of four vehicle types, along with background noise, collected from March 2023 to October 2024. Data were gathered across six distinct real-world scenes, each featuring diverse terrain types (asphalt, dirt, gravel, and concrete) and varying weather conditions (sunny, rainy, and windy).
Each scene is instrumented with a spatially distributed sensor network composed of 6 to 8 nodes. Every node includes a co-located microphone (sampled at 16 kHz) and a geophone (sampled at 200 Hz). Vehicle GPS trajectories were recorded at a rate of 1 Hz. All recordings are segmented into non-overlapping 2-second clips, resulting in a total of 21,694 samples. These clips are transformed into mel-scale spectrograms for model input. GPS coordinates are converted into meter-level spatial positions using the Local Tangent Plane approximation (united1987department).
@@
Multi-Vehicle Joint Classification and Localization. For multi-vehicle settings, we use scene I22, which contains multiple moving vehicles. A two-layer transformer serves as the task head, trained with a DETR-style loss function (carion2020end) to handle set-based predictions.
@@ E.3 Settings for Ridgecrest Seismicity Dataset
This dataset contains seismic waveform recordings from 31,452 earthquake events (M > -0.5) occurring between January 1, 2020, and December 31, 2024, within an 80 km radius of (35.9°, -117.6°) in the Ridgecrest region of California. The data collection and processing procedures largely follow the methodology outlined by Si et al. (si2024all).
We obtained the earthquake event catalog by querying the Southern California Seismic Network (SCSN) (caltech1926socal) via the Southern California Earthquake Data Center (SCEDC) (center2013southern) online catalog. The selected events include magnitudes higher than -0.5 and depths larger than -5 km. For each event, we collected three-component (East, North, Vertical) waveform data from 16 stations in the California Integrated Seismic Network, using two modalities: high-gain broadband seismometers and high-gain accelerometers. All data are sampled at 100 Hz and retrieved in miniSEED format from the SCEDC Open Data repository.
For each event, we extract a 30.72-second window from all channels as model input. During preprocessing, we detrend the waveforms and apply the Short-Time Fourier Transform (STFT) to generate spectrograms. A 2-35-Hz band-pass filter is applied to remove low-frequency noise (e.g., oceanic and atmospheric microseisms) and high-frequency instrumental or environmental noise. We convert spatial positions of each station from GPS signals to kilometer-level positions using Local Tangent Plane projection (united1987department).
@@ E.4 Settings for RealWorld-HAR Dataset
The RealWorld Human Activity Recognition (HAR) dataset (sztyler2016body) comprises multi-modal activity signals collected from 15 participants. The dataset captures eight common activity types: walking, sitting, lying, climbing down, running, standing, climbing up, and jumping.